OpenStudy (anonymous):

Looking for some hints on where to go with an integration problem. I'll post the question and my progress in the comments so I can format things properly.

OpenStudy (anonymous):

Question: For any given function f, let \[I = \int\limits_{}^{}[f'(x)]^2[f(x)]^ndx\]where n is a positive integer. Show tht, if f(x) satisfies \[f''(x) = kf(x)f'(x)\] for some constant k, then I can be evaluated in terms of f(x), f'(x), k and n.

OpenStudy (anonymous):

From the second condition:\[f(x) = \frac{f''(x)}{kf'(x)}\]\[\int\limits f(x) dx = \frac{1}{k}lnf'(x) + c\]Doesn't immediately appear to be useful.

OpenStudy (anonymous):

In the case n=1: \[I = \int\limits f'(x)[f'(x)f(x)]dx = \int\limits f'(x)\frac{f''(x)}{k}dx = \frac{1}{k}\int\limits f'(x)f''(x)dx\]\[= \frac{1}{k}[f'(x)]^2 - \frac{1}{k}\int\limits f'(x)f''(x)dx\]\[2I = \frac{1}{k}[f'(x)]^2\]\[I = \frac{1}{2k}[f'(x)]^2\]I also tried to consider the case n=2 but I wasn't able to get it into a nice form like this.

OpenStudy (ikram002p):

i think we need ordinary differential equation humm ill see if it cud work.

OpenStudy (anonymous):

I also tried integrating straight off the bat by substitution but I wasn't able to get down to an integral that I could evaluate.

ganeshie8 (ganeshie8):

sub \(f'(x) = t\) then, \(f(x) = t^2/2 + C\) \(f''(x) dx = dt \)

ganeshie8 (ganeshie8):

not really sure, hey ikram help

ganeshie8 (ganeshie8):

substitution seems to work, but im not really sure yet... humm

OpenStudy (anonymous):

Will see if I can do something with those suggestions!

ganeshie8 (ganeshie8):

im really not sure cuz ikram said something about partials :o

ganeshie8 (ganeshie8):

*differential equations

OpenStudy (ikram002p):

humm ordinary , nt partial in ordinary u assume f =1 or C f' =x f" = x^2 and so i never used it in case like this :o

OpenStudy (anonymous):

I get as far as \[I = \frac{1}{k}\int\limits t(\frac{t^2}{2}+C)^{n-1}dt\]I was considering looking at some cases of this but it seems to me like the (t^2/2 + C)^(n-1) part is still going to be an issue.

ganeshie8 (ganeshie8):

why ? sub again : \(\large \frac{t^2}{2} +C = u \)

ganeshie8 (ganeshie8):

ikram, do u see any goofup in the substitution in my first reply ?

OpenStudy (ikram002p):

no it works too :)

ganeshie8 (ganeshie8):

oww ty :))

OpenStudy (ikram002p):

np

OpenStudy (anonymous):

So u = f'(x) as well?

OpenStudy (ikram002p):

the integral u reached is also work better :o

ganeshie8 (ganeshie8):

\(I = \frac{1}{k}\int\limits t(\frac{t^2}{2}+C)^{n-1}dt\) yes, this is a good integral

OpenStudy (anonymous):

\[I = \frac{1}{k}\int\limits \frac{du}{dt}u^{n-1}dt = \frac{1}{k}\int\limits u^{n-1}du = \frac{1}{nk}u^n + c = \frac{1}{nk}[f(x)]^n + c\]How does that look to you guys? I'm slightly concerned about the absence of any instances of f'(x); did I make a mistake evaluating somewhere?

OpenStudy (anonymous):

Still doesn't seem quite right to me, any thoughts @ikram002p @ganeshie8?

ganeshie8 (ganeshie8):

\(I = \frac{1}{k}\int\limits t(\frac{t^2}{2}+C)^{n-1}dt\) we can clearly work it when n = 1, for n > 1 u can substitute t^2/2 + c = u, then tdt = du \(I = \frac{1}{k}\int\limits u^{n-1}du\)

ganeshie8 (ganeshie8):

btw, u = t^2/2 + c, its not f(x) u may differentiate it back and see if u get back to the original function

OpenStudy (anonymous):

\[I = \frac{1}{nk}(\frac{f'(x)^2}{2}+c_1)^n + c_2\]Anything we can do with this?